THE PFI-ZEKE PHOTOELECTRON SPECTROSCOPY OF C_{2}H_{2} AND C_{2}D_{2} NEAR ITS IONIZATION THRESHOLD

SHEUNN-JIUN TANG, JIM JER-MIN LIN, AND \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P.O. Box 23-166, Taipei 107, Taiwan, Republic of China, and Department of Chemistry, National Taiwan University, Taipei, Taiwan, R.O.C..

The 1+1’ two-color pulsed-field induced zero-kinetic-energy (PFI-ZEKE) photoelectron spectra of C_{2}H_{2} via a single rovibronic level of the Ã^{1}A_{u} state of acetylene were previously reported by Pratt et al ^{a}. In this work, with similar approach the effects of the pulsed-field strengths and the rovibronic symmetries of the intermediate states of C_{2}H_{2} and C_{2}D_{2} on the PFI-ZEKE photoelectron spectra near the ionization threshold were investigated and will be reported. Although the number of observed transitions was limited, the spectroscopic constants of the zero-point energy levels of C_{2}H_{2}^{+} obtained from the best fit are in good agreement with those reported by Jagod et al ^{b}: B=1.098± 0.002 cm^{−1}, A_{so}(spin-orbit constant)=-30.84±0.05 cm^{−1}, and IP_{0}=91967.06±0.03 cm^{−1}. Those of C_{2}D_{2}^{+} are B=0.82± 0.03 cm^{−1}, A_{so}=-31.1±0.3 cm^{−1}, and IP_{0}=92004.7± 0.3 cm^{−1}. IP_{0} is defined as the energy difference between the zero-point levels of the ground states of the neutral and its cation. We propose to use IP_{0} to denote the ionization potential of the molecule, of which the cation belongs to the Hund’s case (b) coupling.

aS.T. Pratt, P. M. Dehmer, and J. L. Dehmer, J. Chem. Phys. 99, 6233(1993).

b.M.-F. Jagod, M. Rösslein, C. M. Gabrys, B. D. Rehfuss, F. Scappini, M. W. Crofton, and T. Oka, J. Chem. Phys. 97, 7111(1992).